OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $3.36, and a college-ruled notebook for $4.57. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $205.51. How many of each type were sold?

OpenStudy (amistre64):

my formula is: b@y = (z - tx)/(y-x)

OpenStudy (amistre64):

graphs @ 3.36 = (205.51 - 50(4.57))/(3.36-4.57) might have to chop off a negative sign tho

OpenStudy (amistre64):

google calculator says 19 :) so the rest would be colleged

OpenStudy (anonymous):

sorry im confused

OpenStudy (amistre64):

you need to be more specific; what are you confused about?

OpenStudy (amistre64):

the normal long way to do this is to define the seperate amounts that total up to 50 a + b = 50 now write an equation that totals up the number of items times their prices that equal 205.51 ax + by = 205.51 ; and solve either by substitution or elimination i eliminate.................. a + b = 50 <-- times by -x ax + by = 205.51 -ax -bx = -50x ax+by = 205.51 ---------------- b(y-x) = (205.51 - 50x) b@y = (205.51 - 50x)/(y-x)

OpenStudy (amistre64):

determine what you want as b and what its price is; and fill in the restq

OpenStudy (anonymous):

Ummm..amistre? That would be all well and good if the prices weren't already set. I think you're nuking this a bit. Let a = number of graph-paper notebooks sold. Let b = number of college-ruled notebooks sold. a + b = 50 3.36a + 4.57b = 205.51 Solve the system... a = 50-b Substitue into 2nd equation 3.36(50-b) + 4.57b = 205.51 168.00 - 3.36b + 4.57b = 205.51 1.21b = 37.51 b = 31 So, a = 19. 19 graph-paper notebooks were sold and 31 college-ruled notebooks were sold.

OpenStudy (anonymous):

ty makes more sense mtbender:)

OpenStudy (amistre64):

i kept a and b; as well as x and y generic so that it wouldnt matter in the end. if b = graphs at y = 3.36; then a = college at x = 4.57 or if b = college at y = 4.57; then a = graphs at x = 3.36

OpenStudy (anonymous):

I kept up with you, but given the nature of the question, I thought yoyokris wasn't tracking :) I did like your generic formula...took me back to Business Calculus :)

OpenStudy (amistre64):

yay!! :)

OpenStudy (amistre64):

its hard to tell what yoyok is or is not tracking since the only response is "im confused" ...